OpenStudy (lily):

One Monday,the bus carried 364 passengers and 157 were women.What percentage were women?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (robtobey2):

(157/364)100

OpenStudy (lily):

thank you!

OpenStudy (skullpatrol):

Actually (157/364)100%

OpenStudy (lily):

thanks that really helpful!

OpenStudy (skullpatrol):

Thanks for asking :-)

OpenStudy (skullpatrol):

Remember $$\large \text{%} = \dfrac{1}{100}$$

OpenStudy (lily):

Ok I wrote everything down :)

OpenStudy (skullpatrol):

Here is the general formula: $$\large \dfrac{x}{100} = \dfrac{a}{b}$$

OpenStudy (skullpatrol):

The left hand side is just x/100 = x%